# Incremental Model Order Reduction of Smoothed-Particle Hydrodynamic Simulations

### v1

@@ 6 Validation
In line with Eqn. (29), we build an error matrix $E\in\mathbb{R}^{5\times 5000}$, where each row is assigned to a field and each column to the result of a time step. This matrix’s mean and maximum values are indicated by $\varepsilon_{mean}$ and $\varepsilon_{max}$, respectively. Table 3 shows the final rank $q_{\phi}$ of a field $\phi$, the sum of all field ranks $\sum q$, the error measures $\varepsilon_{mean}$ and $\varepsilon_{max}$, as well as the relative computational overhead $t^{SVD}/t^{SPH}\cdot 100\%$ for all utilized imputation strategies. Therein, $t^{SVD}$ and $t^{SPH}$ refer to the measured wall clock time of the ROM construction and the flow solver integration, respectively. The simulations are conducted thrice to underline the credibility of the measured wall clock time, and their mean value is reported. Investigations of this first step are performed for a fixed bunch width of $b=250$ and no windowing was employed. While the different imputation strategies do not significantly affect the accuracy, the associated number of modes and computational effort vary significantly. As expected, the GPOD with 10 iterations achieves the best compression and requires the least number of modes. On the other hand, this example has a CPU overhead that is almost as expensive as the SPH simulation. The BMI method is algorithmically very simple and requires only a moderate computational effort in the range of 30%. However, it still delivers competitive compression rates. The strategy can be further optimized slightly in combination with the hybrid $h_{MI-BMI}$. Hence, the strategy we employ for the remaining investigations is the hybrid $h_{MI-BMI}$ approach.
@@
In a second step, the windowing strategy is assessed for the $h_{MI-BMI}$. The number of particles included in a given SVD and the number of time steps assigned to the SVD are influencing parameters on the compression ratio (cf. Eqn. 31) and the CPU effort. However, since we also store values for inactive particles, the compression ratio (CR) is measured against a Full-Order Full-Storage (FOFS) approach. For each time step ($n_{t}$), one needs to store one numerical value per field ($n_{f}$) per active particle ($n_{p}$), allowing for a hypothetical FOFS approach’s storing effort, proportional to approximately $c\,\mathrm{n_{f}}\,\sum_{i=1}^{i=n_{t}}\,\mathrm{n_{p}(i)}$, where $c$ denotes the required size for storing a number, that depends, e.g., on the underlying single- or double precision metric. Table 4 presents the results for three different windowing configurations and four different bunch-widths. The compression ratio moderately varies between $7.72\leq CR\leq 9.03$ and reveals minor benefits for the single window approach. The relative computational overhead $t^{SVD}/t^{SPH}\cdot 100\%$ is provided in Tab. 5. In contrast to the compression ratio, the computational overhead varies substantially, between $7.6\leq t^{SVD}/t^{SPH}\cdot 100\%\leq 50.0$. The overhead lowers as the number of windows increases for all considered bunch-widths. Additionally, a larger number of windows seems to favor smaller bunch widths, as minimal efforts are continuously shifted towards smaller bunch-widths when the amount of windows increases.
@@
Regarding SVD, the case shares many properties with the example discussed in Sec. 6.2, particularly the presence of input and output boundaries where particles can enter and exit the domain. However, unlike the previous scenario, no fully developed state is reached where the number of active liquid particles remains virtually constant. As a result, the reduced system exhibits a fully unsteady behavior. The 3D test case is also significantly more computationally intensive and is therefore computed using a thread-parallel approach on one GPU.
@@
With regard to pressure, however, another detail comes to the fore, which concerns the solid pressures and their reconstruction. The solid pressures in SPH are usually determined from the fluid properties using boundary-specific formulae and the algebraic equation of state (5). The pressure levels can quickly be in the range of several million Pascals in wetted regions, but fall back to zero Pascal in (temporarily) unwetted solid areas. The reconstruction of the pressure field by the SVD is therefore more challenging for the solid than in the fluid particles. In essence, the snapshot matrix for the pressure of a wall particle resembles a random matrix with many zero entries, and even when maintaining a large number of modes, it is difficult to maintain the same accuracy as for the fluid particles. Alternatively, analogous to the procedure for the FOM SPH method, the reconstructed fluid pressure can be determined in a post-processing step, which appears to be much simpler. Nonetheless, we obtain the ROM pressure field for the solid from the SVD of the FOM solid pressures in the present study, to evaluate the forces and torques on the turbine. Mind that only a subset of solid particles interacts with water during the simulation; for those that do not interact, the pressure value remains zero. Solid particles are included in their corresponding bunch matrix only if they have interacted with fluid particles at least once. This avoids storing irrelevant data. Solid particles are then handled similarly to newly injected fluid particles, with the distinction that imputation is unnecessary since no data is missing.
@@
Figure 15 depicts the maximum (left) and mean (right) reconstruction errors obtained for the solid pressure using three different (fixed) truncation ranks in the range of 25%-75% of the maximum possible rank. The irregularity of the snapshot matrix of the solid pressures mainly concerns the maximum error. The mean error, however, remains relatively small, which suggests that the rank truncation primarily impairs the local accuracy of the wall pressure. The latter is confirmed by evaluating the reconstruction of integral quantities, i.e., forces and torques, as shown in Figs. 16-17, revealing an error is in the order of per mille.
@@
The attained compression ratio (CR) for the Pelton runner case is evaluated against the full-order full-storage effort in Tab. 6. This assessment is performed separately for the fluid and solid particles and subsequently collectively. The tabulated values for the fluid data are of the same order of magnitude as those in the 2D cases, cf. for example Tab. 4, and the experienced CR reduction can be attributed the ROM of the solid pressure. Mind that post-processing the ROM for the solid pressures from the ROM fluid data in line with the employed boundary condition could represent a more efficient strategy. The computational overhead is $t^{SVD}/t^{SPH}\approx 7.26\%$ for the considered simulation which agrees with best practice results obtained for the 2D impinging jet case depicted in Tab 5.

### v2

@@ 6 Validation
(a)   FOM time step = 250
(b)   ROM time step = 250
(c)   FOM time step = 2500
(d)   ROM time step = 2500
(e)   FOM time step = 5000
(f)   ROM time step = 5000
Figure 13: 2D Impinging Jet: Snap-shot of the particle positions colored by dynamic pressure field obtained from the original FOM (left; SPH) and the SVD-based reconstruction (right) at three different time instants.
In line with Eqn. (31), we build an error matrix $E\in\mathbb{R}^{5\times 5000}$, where each row is assigned to one of the five physical fields stored — here in particular the two components of position, the two components of velocity, and the density — and each column corresponds to a particular time step. This matrix’s mean and maximum values are indicated by $\varepsilon_{mean}$ and $\varepsilon_{max}$, respectively. Table 3 shows the final rank $q_{\phi}$ of a field $\phi$, the sum of all field ranks $\sum q$, the error measures $\varepsilon_{mean}$ and $\varepsilon_{max}$, as well as the relative computational overhead $t^{SVD}/t^{SPH}\cdot 100\%$ for all utilized imputation strategies. Therein, $t^{SVD}$ and $t^{SPH}$ refer to the measured wall clock time of the ROM construction and the flow solver integration, respectively. The simulations are conducted thrice to underline the credibility of the measured wall clock time, and their mean value is reported. Investigations of this first step are performed for a fixed bunch width of $b=250$ and no windowing was employed. While the different imputation strategies do not significantly affect the accuracy, the associated number of modes and computational effort vary significantly. Considering that the imputation strategies were tested on the same simulation, the number of retained modes directly relates to the compression rate. As expected, the GPOD with 10 iterations achieves the best compression and requires the least number of modes. On the other hand, this example has a CPU overhead that is almost as expensive as the SPH simulation. The BMI method is algorithmically very simple and requires only a moderate computational effort in the range of 30%. However, it still delivers competitive compression rates. The strategy can be further optimized slightly in combination with the hybrid $h_{MI-BMI}$. Hence, the strategy we employ for the remaining investigations is the hybrid $h_{MI-BMI}$ approach.
@@
In a second step, the windowing strategy is assessed for the $h_{MI-BMI}$. The number of particles included in a given SVD and the number of time steps assigned to the SVD are influencing parameters on the compression ratio (cf. Eqn. 33) and the CPU effort. However, since we also store values for inactive particles, the compression ratio (CR) is measured against a Full-Order Full-Storage (FOFS) approach. For each time step ($n_{t}$), one needs to store one numerical value per field ($n_{f}$) per active particle ($n_{p}$), allowing for a hypothetical FOFS approach’s storing effort, proportional to approximately $c\,\mathrm{n_{f}}\,\sum_{i=1}^{i=n_{t}}\,\mathrm{n_{p}(i)}$, where $c$ denotes the required size for storing a number, that depends, e.g., on the underlying single- or double precision metric. Table 4 presents the results for three different windowing configurations and four different bunch-widths. The compression ratio moderately varies between $7.72\leq CR\leq 9.03$ and reveals minor benefits for the single window approach. The relative computational overhead $t^{SVD}/t^{SPH}\cdot 100\%$ is provided in Table 5. In contrast to the compression ratio, the computational overhead varies substantially, between $7.6\leq t^{SVD}/t^{SPH}\cdot 100\%\leq 50.0$. The overhead lowers as the number of windows increases for all considered bunch-widths. Additionally, a larger number of windows seems to favor smaller bunch widths, as minimal efforts are continuously shifted towards smaller bunch-widths when the amount of windows increases.
@@
Regarding SVD, the case shares many properties with the example discussed in Sec. 6.2, particularly the presence of input and output boundaries where particles can enter and exit the domain. However, unlike the previous scenario, no fully developed state is reached where the number of active liquid particles remains virtually constant, but the total number of simultaneously active particles at the end of the simulation grows to over 620 000. As a result, the reduced system exhibits a fully unsteady behavior. The 3D test case is also significantly more computationally intensive and is therefore computed using a thread-parallel approach on one GPU.
@@
With regard to pressure, however, another detail comes to the fore, which concerns the solid pressures and their reconstruction. The solid pressures in SPH are usually determined from the fluid properties using boundary-specific formulae and the algebraic equation of state (5). The pressure levels can quickly be in the range of several million Pascals in wetted regions, but fall back to zero Pascal in (temporarily) unwetted solid areas. The reconstruction of the pressure field by the SVD is therefore more challenging for the solid than in the fluid particles. In essence, the snapshot matrix for the pressure of a wall particle resembles a random matrix with many zero entries, and even when maintaining a large number of modes, it is difficult to maintain the same accuracy as for the fluid particles. For this reason, the adaptive truncation approach would lead to storing almost all the available modes. Therefore, we decided to use a fixed truncation rank for the pressure field, which allows assigning the degree of compression but not the resulting accuracy. The accuracy of the reconstruction is evaluated based on the maximum and mean error, as well as for the integral quantities force and moment. Alternatively, analogous to the procedure for the FOM SPH method, the reconstructed fluid pressure can be determined in a post-processing step, which appears to be much simpler. Nonetheless, we obtain the ROM pressure field for the solid from the SVD of the FOM solid pressures in the present study, to evaluate the forces and torques on the turbine. Mind that only a subset of solid particles interacts with water during the simulation; for those that do not interact, the pressure value remains zero. Solid particles are included in their corresponding bunch matrix only if they have interacted with fluid particles at least once. This avoids storing irrelevant data. Solid particles are then handled similarly to newly injected fluid particles, with the distinction that imputation is unnecessary since no data is missing.
@@
Figure 17 depicts the maximum (left) and mean (right) reconstruction errors obtained for the solid pressure using three different (fixed) truncation ranks in the range of 25%-75% of the maximum possible rank. The irregularity of the snapshot matrix of the solid pressures mainly concerns the maximum error. The mean error, however, remains relatively small, which suggests that the rank truncation primarily impairs the local accuracy of the wall pressure. The latter is confirmed by evaluating the reconstruction of integral quantities, i.e., forces and torques, as shown in Figs. 18-19, revealing an error is in the order of per mille.
@@
The attained compression ratio (CR) for the Pelton runner case is evaluated against the full-order full-storage effort in Table 6. This assessment is performed separately for the fluid and solid particles and subsequently collectively. The tabulated values for the fluid data are of the same order of magnitude as those in the 2D cases, cf. for example Table 4, and the experienced CR reduction can be attributed the ROM of the solid pressure. Mind that post-processing the ROM for the solid pressures from the ROM fluid data in line with the employed boundary condition could represent a more efficient strategy. The computational overhead is $t^{SVD}/t^{SPH}\approx 7.26\%$ for the considered simulation which agrees with best practice results obtained for the 2D impinging jet case depicted in Table 5.
@@ 8 CRediT Authorship Contribution Statement
9 Data Availability Statement
Data sharing not applicable, since no new data was generated, but only standard 2D testcases were used to scrutinze the method. For the final 3D testcase, the employed geometry is confidential, but the method essentially works for any irregular data matrix.